Convert into Fahrenheit?

1.) 34.1 Celsius

2.) 28.3 Celsius

3.) 36.8 Celsius

4.) 41.2 Celsius

5.) 37.7 Celsius
Answer #1

Ans) To convert Frahenheit to Celsius:

Formula: C= 5/9(F-32)

But here in question they are asking to convert Celsius into Frahenheit

Formula: F= (C×9/5) + 32

1) 34.1°C to °F

Solution:

(34.1°C × 9/5) + 32 = 93.38°F

2) 28.3°C to °F

Solution:

(28.3°C × 9/5) + 32 = 82.94°F

3) 36.8°C to °F

Solution:

(36.8°C × 9/5) + 32 = 98.24°F

4) 41.2°C to °F

Solution:

(41.2°C × 9/5) + 32 = 106.16°F

5) 37.7°C to °F

Solution:

(37.7°C × 9/5) + 32 = 99.86°F